OpenStudy (anonymous):

What value of c makes the polynomial below a perfect square? x^2+10x+c C=?

jimthompson5910 (jim_thompson5910):

the x coefficient is 10 take half of this to get 10/2 = 5 then square this to get 5^2 = 25

jimthompson5910 (jim_thompson5910):

so c = 25

OpenStudy (anonymous):

Thank you so much for helping me !!((: i appreciate it

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

what about : what value of b makes the polynomial below a perfect square? x^2+bx+16 B=??

jimthompson5910 (jim_thompson5910):

go backwards from what I did

jimthompson5910 (jim_thompson5910):

you first take the square root of 16 to get sqrt(16) = 4 then you double this to get 4*2 = 8 that means b = 8

jimthompson5910 (jim_thompson5910):

notice how I undid everything I did up top

OpenStudy (anonymous):

Your like super good at math, i wish i wasss :c !!

jimthompson5910 (jim_thompson5910):

keep practicing and you'll get better

OpenStudy (anonymous):

But yes i understand what your sayingg

jimthompson5910 (jim_thompson5910):

that's good, and that's I how got to be so good at math: lots and lots of practice

OpenStudy (anonymous):

Ive been practicing .. math is the hardest subject for me i was never good at it

jimthompson5910 (jim_thompson5910):

well try to translate everything to what you can connect to that usually helps me

OpenStudy (anonymous):

Thanks for the advice c:

jimthompson5910 (jim_thompson5910):

you're welcome